OpenStudy (anonymous):

What is the minimum number of bits required to represent +56,392 using 2’s complement form? conversion: Decimal to binary... grrr.. : / need help quick...

OpenStudy (anonymous):

signed numbers excluding 0 = 2^5-1 digits including 0 = 2^5 digits

OpenStudy (anonymous):

i got it thanks :)

OpenStudy (anonymous):

just for binary, the number if bits is \[\lceil \log_{2}(56392) \rceil = 16\] For 2's compliment, add 1. (This matches the well known limits for 16-bit signed/unsigned values; maximum 16-bit unsigned = 2¹⁶-1 = 65535, maximum 16-bit signed = 2¹⁵-1 = 32767. 56392 can be expressed in a 16-bit unsigned value, but not in a 16-bit signed value.)